When simplified and expressed with negative exponents,


the expression (x + y)-1 (x-1 + y-1) is equal to:


(A ) x-2 + 2x-ly-l + y-2 (B ) x-2 + 2- lx-ly-1 + y-2


(C) x-1 y-1 (D ) x-2 + y-2

Grade:Upto college level

1 Answers

Pratik Tibrewal
10 years ago
(x+y)^-1 (1/x + 1/y) = 1/(x + y) * ( x+y)/xy = 1/xy = x^-1 y^-1
(C)


Thanks and Regards,
Pratik Tibrewal
askiitians faculty
BTech